### v1

# Unraveling the Connections between Flow Matching and Diffusion Probabilistic Models in Training-free Conditional Generation

Abstract
Training-free conditional generation aims to leverage the unconditional diffusion models to implement the conditional generation, where flow-matching (FM) and diffusion probabilistic models (DPMs) are two mature unconditional diffusion models that achieve high-quality generation. Two questions were asked in this paper: What are the underlying connections between FM and DPMs in training-free conditional generation? Can we leverage DPMs to improve the training-free conditional generation for FM? We first show that a probabilistic diffusion path can be associated with the FM and DPMs. Then, we reformulate the ordinary differential equation (ODE) of FM based on the score function of DPMs, and thus, the conditions in FM can be incorporated as those in DPMs. Finally, we propose two posterior sampling methods to estimate the conditional term and achieve a training-free conditional generation of FM. Experimental results show that our proposed method could be implemented for various conditional generation tasks. Our method can generate higher-quality results than the state-of-the-art methods.
Figure 1: Our proposed FMPS in pre-trained pixel and latent flow diffusion models in various conditional generation applications. FMPS has the same flexibility and general as the training-free methods for diffusion probabilistic models.
1 Introduction
Diffusion-based conditional generation methods [1, 2, 3] aims at generating desired images according to the given conditions, such as text prompts. Conditional generation is the key component for various computer vision tasks, such as linear/non-linear inverse problems [2, 1], image edit [4], etc. These methods can be roughly categorized into training-based [5, 6] and training-free [7, 3] approaches. In this paper, we mainly focus on the training-free method.
Training-free conditional generation methods [1, 3, 8] rely on a pre-trained unconditional diffusion model and do not require retraining. Based on the way to build the probabilistic path from Gaussian noise to data distribution, the unconditional diffusion models can be divided into two categories: diffusion probabilistic models (DPMs) [9, 10, 11], and flow diffusion models [12]. DPMs introduced the score function [9] to model the probabilistic path as a stochastic path. Flow diffusion models used Continuous Normalizing Flow (CNF) [13] to model the probabilistic path via flow-matching (FM), where CNF enables an interpolation way to define the probabilistic path from two distribution points. To simplify the notation, we abbreviated flow diffusion models as FMs to emphasize the flow-match.
When the unconditional diffusion model is a DPM, one of the training-free conditional DPMs is based on the posterior sampling [2]. DPS [2] first proposed the posterior sampling to solve the linear inverse problems. FreeDom [2] improved DPS by introducing the energy-based function to extend posterior sampling to solve non-linear inverse problems. MPGD [3] further introduced the manifold theory to improve FreeDom. The posterior sampling methods provide a flexible solution for various conditional generation tasks.
Unlike DPMs, there was very little research on training-free conditional FMs. The latest works include D-flow [14] and $\pi$-flow [15]. D-flow guided the generation process by fine-tuning the start point of FMs. $\pi$-flow [15] only focuses on the linear inverse problems for conditional generation.
The existing training-free methods for FMs have certain limitations, e.g., the $\pi$-flow is hidden on the linear inverse assumption. While the posterior sampling [2] in DPMs can incorporate the conditions into the unconditional diffusion model easily. Unlike DPMs, FM does not have a direct way of introducing the conditions. Therefore, two natural questions arise: 1) Is there any connections between FM and DPMs? and 2) Can we leverage posterior sampling in DPMs to improve the training-free conditional generation for FM?
In this paper, we propose a novel training-free conditional generation method for FMs called flow-match based posterior sampling (FMPS). The key insight is first to unravel the connection, where there is a probabilistic path between FMs and DPMs. This enables FMPS to define the score function explicitly. In this condition, FMPS could achieve the training-free conditional generation following a similar road of DPMs via posterior sampling.
Concretely, posterior sampling depends on the score function in DPMs. We first explicitly connect the ODE of FM and the score function in DPMs (Detailed in Sec. 4). Then, to make the posterior sampling available for FMs, we redefine the ODE of the FMs as the score function in DPMs. This new definition could introduce conditions into FMs via posterior sampling, where FMPS can also easily change the marginal distribution (i.e., the unconditional FMs) to the conditional distribution (i.e., the conditional FMs). Hence, an unconditional FM and a distance metric could be used for the training-free conditional FMs. Moreover, we also propose two posterior sampling methods: gradient-aware FMPS and gradient-free FMPS, where the difference is whether we need to calculate the gradient of the FMs to perform the posterior sampling. The experimental results on linear inverse problems, non-linear inverse problems, and text-to-image generation tasks show that FMPS achieves better generation quality in various downstream tasks and verify the validity and efficiency of FMPS.
To sum up, the main contributions of the FMPS are:
  • We propose a novel training-free conditional generation method, called FMPS, for the flow diffusion models.
  • We unravel the connections between flow-matching and diffusion probabilistic models.
  • The experimental results show that FMPS could solve board-range conditional generation tasks while maintaining high-quality generation.
2 Related Work
Diffusion-based conditional generation tasks. Diffusion-based conditional generation tasks aim at leveraging the potential of the diffusion models [10, 11, 9] to generate the desired samples, which is an important challenge. Conditional generation has been widely applied in various ways. For example, the inverse problem [16, 2] uses degenerated images as a condition to enhance its quality. The text-to image [17, 18, 19, 20], text-to-video [21], and text-to-3D [22] generations use text prompts as a condition to generate the images, the videos, the 3D objects described by the text prompt, respectively. Depending on whether the condition generation model needs to be retrained, the conditional generation methods can be mainly divided into training-based and training-free methods. In this paper, we focus on the training-free methods.
Training-free methods [2, 1, 3, 16, 15] aim at leveraging pre-trained unconditional diffusion models. Since the unconditional diffusion models can be categorized into DPMs and FMs, the training-free conditional DPMs and the training-free conditional FMs are discussed.
Training-free conditional DPMs. These methods [2, 1, 3, 23] use DPMs as the unconditional diffusion models, where DPMs model the sampling process [11, 10, 9] based on the score function [9]. One of the successful approaches is based on the posterior sampling proposed by DPS [2]. DPS indicated that the score function represents the correlation between marginal distribution, i.e., the target of the unconditional generation and the sampling process. Thus, DPS derived a conditional term to connect the unconditional generation and condition, leveraging this correction. Meanwhile, DPS further proposed an estimation method for the conditional term under linear inverse assumption [16]. Then, FreeDom [1] introduced the energy function to estimate the conditional term without the limitation of linear inverse assumption. MPGD [3] and DSG [7] introduced the manifold theory and sphere constraint to enhance the FreeDom, respectively.
Training-free conditional FMs. Recently, FMs [12, 13] have been proposed to leverage an ODE to model the sampling process via flow-matching [12]. Many researches [18, 13] show the potential of FMs, which makes it a potential choice for training-free methods. Since there is no explicit score function, training-free methods for FMs aim at building the connection between the unconditional generation process and condition. $\pi$Flow first proposed a closed-form after introducing the condition into the flow path under the linear inverse assumption. D-Flow [14] fine-tuned the initialized Gaussian noise to build the connection condition and generated images.
However, training-free conditional FMs are not flexible compared to the DPMs. In this paper, we observe an implicit score function hidden in the ODE of FMs. Leveraging this, we explicitly define the score function-based ODE for FMs. After the score function is combined, we could achieve flexible generation via FMs, where the ability of the FMs bootstraps the generation quality.
3 Preliminary
3.1 Training-free conditional DPMs
Unconditional diffusion probabilistic models. It introduces the Gaussian probabilistic path to model the forward process of Eq. 7 as:
$x_{t}=\sqrt{\hat{\alpha}_{t}}x_{0}+(1-\hat{\alpha}_{t})\epsilon,$ (1)
where $\epsilon\sim\mathcal{N}(0,1)$, $\hat{\alpha}_{t}$ is the parameter of the noise scheduler related to the $t$-th time, $x_{0}$ is the image sampled from the data distribution.
Then, the reverse process of DPMs could be defined based on the score function:
$dx_{t}=[f(t)x_{t}-\frac{1}{2}g(t)^{2}\nabla_{x_{t}}\log p(x_{t})]dt,$ (2)
where $f(t)$ and $g(t)$ are also the parameter of noise scheduler related to the $\hat{\alpha}_{t}$. $\nabla_{x_{t}}\log p(x_{t})$ is the score function, where it can also be represented as the predicted noise formulation: $\epsilon_{\theta}(x_{t},t)$ and is modeled by the neural network.
Training-free conditional DPMs via unconditional DPMs. Given a condition $c$ such as a text prompt, we could change the unconditional score function to the condition distribution as:
$\nabla_{x_{t}}\log p(x_{t}|c).$ (3)
Then, by the Bayesian rule, Eq. 3 could be split as:
$\nabla_{x_{t}}\log p(x_{t}|c)=\nabla_{x_{t}}\log p(x_{t})+\nabla_{x_{t}}\log p(c|x_{t}).$ (4)
$\nabla_{x_{t}}\log p(x_{t})$ is exactly the unconditional score function, which can be used directly. But $\nabla_{x_{t}}\log p(c|x_{t})$ is the conditional term, which is unknown and needs to be estimated.
Posterior Sampling on DPMs. Previous works [1, 2] proposed posterior sampling for the estimation of $\nabla_{x_{t}}\log p(c|x_{t})$. In this paper, we follow Freedom [1], which indicates we could use a differentiate distance metric $D(\ast,\ast)$ to measure the distance between $x_{t}$ and $c$. That is, $\nabla_{x_{t}}\log p(c|x_{t})$ could be directly estimated by $D(\ast,\ast)$ as:
$\nabla_{x_{t}}\log p(c|x_{t})\approx r\nabla_{x_{t}}D(\hat{x}_{0|t},c),$ (5)
where $r$ is the hyperparameter to control the strength of given condition $c$. $\hat{x}_{0|t}$ is the data distribution prediction for $x_{t}$, which is defined as:
$\hat{x}_{0|t}=\frac{1}{\hat{a}_{t}}(x_{t}-(1-\hat{a}_{t})\epsilon_{\theta}(x_{t},t)).$ (6)
Therefore, by Eq. 4 - Eq. 6, we could achieve the training-free conditional generation.
3.2 Flow diffusion models
Flow match-based diffusion models (FMs) define the forward process between data distribution $p_{0}$ and the Gaussian noise distribution $p_{1}$ as follows:
$x_{t}=a_{t}x_{0}+b_{t}\epsilon,$ (7)
where $\epsilon\sim p_{1}=\mathcal{N}(0,1)$, $x_{0}\sim p_{0}$ and $x_{t}$ is the intermediate result. $a_{t}$ and $b_{t}$ are schedule parameters [12].
Then, FMs define the reverse process based on the ODE as:
$dx_{t}=v_{\theta}(x_{t},t)dt,$ (8)
where $v_{\theta}(x_{t},t)$ is called the velocity function, which is modeled by the neural networks.
Thus, the reverse process could be calculated by the ODE solver such as Euler solve:
$x_{t-1}=x_{t}+hv_{\theta}(x_{t},t),$ (9)
where $h$ is the time interval between $t$-th time step to $t-1$-th time step.
Motivation. Focusing on Eq. 7 and Eq. 1, we could find that there is an implicit score function hidden in the ODE of FMs. Since we could set $a_{t}=\sqrt{\hat{\alpha}_{t}}$ and $b_{t}=1-\hat{\alpha}_{t}$. Then, the path from the FMs and DPMs are the same. Thus, the ODE of FMs could be represented by the score function similar to the DPMs. If we can explicitly define the score function to re-formulate the ODE of FMs, posterior sampling could help FMs achieve conditional generation again!
4 Methodology
In this paper, we propose a novel training-free conditional generation method for FMs that leverages posterior sampling in DPMs, called flow-matching based posterior sampling (FMPS). The first step for FMPS is to re-formulate the ODE of the reverse process via the score function. This enables FMPS to leverage the Bayesian rule, the key step for posterior sampling, to introduce the condition $c$ by changing the marginal distribution to the conditional distribution. In the end, using our new conditional term-based reverse process, FMPS proposes two new methods to estimate it and achieve conditional generation.
Connections between FMs and DPMs in the training-free conditional generation. To re-formulate the ODE of the reverse process, we have the following proposition:
Proposition 1 ([24]).
Given any ODE in FMs, there exists a probabilistic diffusion path, which could be represented related to the score function in DPMs as follows:
$v_{\theta}(x_{t},t)=\frac{\dot{b}_{t}}{b_{t}}x_{t}-a_{t}(\dot{a}_{t}-\frac{\dot{b}_{t}}{b_{t}}a_{t})\nabla_{x_{t}}\log p(x_{t}),$ (10)
where $\dot{b}_{t}$ is $\frac{\partial b_{t}}{\partial t}$ and $\dot{a}_{t}$ is $\frac{\partial a_{t}}{\partial t}$.
The detailed proof is offered in the supplementary. Proposition 1 indicates one way to explicitly define the score function in the reverse process of the FMs. Thus, we have the following Lemma.
Lemma 1.
By introducing $c$, the ODE of FMs could be re-formulated based on the score function as follows:
$dx_{t}=[\lambda x_{t}-\beta_{t}(\nabla_{x_{t}}\log p(x_{t})+\nabla_{x_{t}}\log p(c|x_{t}))]dt,$ (11)
where $\lambda=\frac{\dot{b}_{t}}{b_{t}}$ and $\beta=a_{t}(\dot{a}_{t}-\frac{\dot{b}_{t}}{b_{t}}a_{t})$
Proof.
By Eq. 8 and Eq. 10, we could get the unconditional FMs formulation via the score function as:
$dx_{t}=[\frac{\dot{b}_{t}}{b_{t}}x_{t}-a_{t}(\dot{a}_{t}-\frac{\dot{b}_{t}}{b_{t}}a_{t})\nabla_{x_{t}}\log p(x_{t})]dt.$ (12)
To simplify Eq. 12, we set $\lambda=\frac{\dot{b}_{t}}{b_{t}}$ and $\beta_{t}=a_{t}(\dot{a}_{t}-\frac{\dot{b}_{t}}{b_{t}}a_{t})$, thus having:
$dx_{t}=[\lambda x_{t}-\beta_{t}\nabla_{x_{t}}\log p(x_{t})]dt.$ (13)
We introduce the $c$ by changing the marginal distribution $p(x_{t})$ to the conditional distribution $p(x_{t}|c)$ following the posterior sampling as follows:
$dx_{t}=[\lambda x_{t}-\beta_{t}\nabla_{x_{t}}\log p(x_{t}|c)]dt.$ (14)
Then, by the Bayes’ rule, we have:
$p(x_{t}|c)=\frac{p(c|x_{t})p(x_{t})}{\int{p(c)}}.$ (15)
We extent Eq. 15 as:
$\begin{split}\nabla_{x_{t}}\log p(x_{t}|c)&=\nabla_{x_{t}}\log\frac{p(c|x_{t})p(x_{t})}{\int{p(c)}}\\ &=\nabla_{x_{t}}[\log p(c|x_{t})+\log p(x_{t})-\log\int{p(c)}].\end{split}$ (16)
Since $\int{p(c)}$ is irrelevant to the $x_{t}$, $\nabla_{x_{t}}\int{p(c)}$ is zero. With Eq. 16 to Eq. 14, we have:
$dx_{t}=[\lambda x_{t}-\beta_{t}(\nabla_{x_{t}}\log p(x_{t})+\nabla_{x_{t}}\log p(c|x_{t}))]dt.$ (17)
Thus, we finish the proof. ∎
Lemma 1 re-defines a new ODE based on the score function for FMs. In this condition, posterior sampling could work on the FMs again, similar to the posterior sampling on the DPMs.
Posterior sampling for training-free conditional FMs. The next step is to calculate Eq. 11. Under the FMs, $\nabla_{x_{t}}\log p(x_{t})$ and $\nabla_{x_{t}}\log p(c|x_{t})$ could not directly calculate. This is different from the DPMs since we know that $\nabla_{x_{t}}\log p(x_{t})=\epsilon_{\theta}(x_{t},t)$, which is the output of the pre-trained backbone in DPMs. But FMs lose this property.
To achieve posterior sampling, we first implement a math trick for Eq. 11 as follows:
$dx_{t}=[(\lambda x_{t}-\beta_{t}\nabla_{x_{t}}\log p(x_{t}))-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt.$ (18)
Then, we could formulate it with a known unconditional term $v(x_{t},t)$ and a conditional term formulation $\nabla_{x_{t}}\log p(c|x_{t})$ by Eq. 10 as follows:
$dx_{t}=[v_{\theta}(x_{t},t)-\beta_{t}\nabla_{x_{t}}\log p(c|x_{t})]dt.$ (19)
Since $\nabla_{x_{t}}\log p(c|x_{t})$ derives from the probabilistic path, it has the same property and formulation on the DPMs. Following the same estimation from training free conditional DPMs in Eq. 5, we have:
$\nabla_{x_{t}}\log p(c|x_{t})\approx r\nabla_{x_{t}}D(\hat{x}_{0|t},c).$ (20)
Finally, we have to estimate $\hat{x}_{0|t}$ based on the $x_{t}$ via FMs. FMPS offers gradient-aware FMPS (FMPS-gradient) and gradient-free FMPS (FMPS-free) versions by whether $\nabla_{x_{t}}\log p(c|\hat{x}_{0|t})$ needs to introduce the gradient of the diffusion model.
The FMPS-gradient derives from the one-step Euler solver, which could be represented as follows:
$\hat{x}_{0|t}=x_{t}-tv_{\theta}(x_{t},t).$ (21)
Eq. 21 directly solves the $t\rightarrow 0$ interval to finish the estimation for $x_{t}$. The advantage of gradient-aware estimation is introducing the gradient information of $\frac{\partial v(x_{t},t)}{\partial x_{t}}$ when calculating $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. Gradient information will benefit the conditional tasks [25]. However, it will introduce additional computation costs.
This motivates us to propose our FMPS-free. We re-define the forward process of FMs as follows:
$x_{0}=\frac{1}{a_{t}}(x_{t}-b_{t}x_{1}).$ (22)
Eq. 22 shows $x_{0}$ is related to the intermediate state $x_{t}$ and the initial state $x_{1}$. Since $x_{1}$ is known and unchanged during the overall generation process and $x_{t}$ is already calculated in the $t$-th time step, we could directly use Eq. 22 to estimate $\hat{x}_{0|t}$ as:
$\hat{x}_{0|t}=\frac{1}{a_{t}}(x_{t}-b_{t}x_{1}).$ (23)
Eq. 23 decouples the connect between $\hat{x}_{0|t}$ and $v_{\theta}(x_{t},t)$ compared to Eq. 21. Thus, it avoids calculate the gradient $\frac{\partial v(x_{t},t)}{\partial x_{t}}$ when calculating $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$, which achieves the gradient-free estimation. This enables saving a lot of the computation costs but needs to skip the first step since when $t=1$, $a_{t}=0$.
Normalization trick for FMPS. According to the DPMs [25, 7], there is an unstable phenomenon: $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will be over powerful or too weaken compared to the unconditional term, which leading the generated direction to deviate from the data distribution and condition, respectively. To alleviate this, we propose a normalization trick to limit the norm of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. Concretely, ODE decides the strength of $v(\ast,\ast)$ could be measured by its norm since $v(\ast,\ast)$ represents a forward direction in ODE. Thus, we could measure the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$. We have:
$g^{1}=\frac{||v(x_{t},t)||_{2}}{||\nabla_{x_{t}}D(\hat{x}_{0|t},c)||_{2}}\nabla_{x_{t}}D(\hat{x}_{0|t},c).$ (24)
Eq. 24 limits the strength of $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ under the bound of $||v(x_{t},t)||_{2}$, where $\nabla_{x_{t}}D(\hat{x}_{0|t},c)$ will not too strong and too weaken, thus alleviating the unstable phenomenon.
Algorithm 1 The algorithm for FMPS-gradient
1:Input: $v_{\theta}(\ast,\ast)$, $T$, $c$, $D(\ast,\ast)$, $r$
2:Output: $x_{0}$
3:Initialize time interval $n\leftarrow\frac{1}{T}$
4:$x_{T}\sim\mathcal{N}(0,1)$
5:$x_{t}\leftarrow x_{T}$
6:for $t$, $t-1$ in pair $[(1,1-n),(1-n,1-2n),(1-in,1-(i+1)n)...,(1-(T-1)n,0)]$ do
7:     Calculate $a_{t}$, $b_{t}$, $\dot{a}_{t}$ and $\dot{b}_{t}$ by Eq. 7.
8:     Calculate $\lambda_{t}$ and $\beta_{t}$ by Eq. 11.
9:     Calculate $v_{\theta}(x_{t},t)$.
10:     Calculate $\hat{x}_{0|t}$ by Eq. 21.
11:     Calculate $g^{1}$ by Eq. 24.
12:     Calculate $x_{t-1}$ using Euler solver by Eq. 26.
13:     $x_{t}\leftarrow x_{t-1}$.
14:end for
15:Return: $x_{0}$
Thus, we could get the conditional generation ODE for FMPS as follows:
$dx_{t}=[v_{\theta}(x_{t},t)-r\beta_{t}g^{1}]dt.$ (25)
The related Euler solve for Eq. 25 is:
$x_{t-1}=x_{t}+h(v_{\theta}(x_{t},t)-r\beta_{t}g^{1}).$ (26)
The overall framework for FMPS based on gradient-aware and gradient-free estimations for $\hat{x}_{0|t}$ are summarized as Algorithm 1 and Algorithm 2, respectively. To sum up, FMPS first leverages the score function in DPMs to re-formulate the ODE of FMs. The score function enables us to introduce the condition directly by changing the marginal distribution to the conditional distribution. In this way, posterior sampling could work for FMs. Thus, we can propose a flow-based posterior sampling to achieve conditional generation.
Additionally, we further introduce a normalization trick to enhance FMPS and improve generation quality. Besides, FMPS also offers a gradient-free estimation for $\hat{x}_{0|t}$, which provides an alternative method to reduce the time costs for calculating posterior sampling.
Algorithm 2 The algorithm for FMPS-free
1:Input: $v_{\theta}(\ast,\ast)$, $T$, $c$, $D(\ast,\ast)$, $r$
2:Output: $x_{0}$
3:Initialize $n$, $x_{T}$, and $x_{t}$. $\triangleright$ Algorithm 1.
[…]
5:for $t$, $t-1$ in pair $[(1,1-n),(1-in,1-(i+1)n),...,(1-(T-1)n,0)]$ do
6:     Calculate $a_{t}$, $b_{t}$, $\dot{a}_{t}$, $\dot{b}_{t}$, $\lambda_{t}$, $v_{\theta}(x_{t},t)$ and $\beta_{t}$. $\triangleright$ Algorithm 1.
7:     if Isfirst then
8:         Calculate $x_{t-1}$ by Eq. 9 $\triangleright$ Skip first step.
[…]
10:     else
11:         Calculate $\hat{x}_{0|t}$ by Eq. 23.
12:         Calculate $x_{t-1}$triangleright Algorithm 1.
13:     end if
14:     $x_{t}\leftarrow x_{t-1}$.
15:end for
16:Return: $x_{0}$
Table 1: Qualitative evaluation of linear inverse problems on CelebA-HQ, where SR (4X) represents the super-resolution (4X), and Deblur represents Gaussian Deblur.
Method Inpainting SR (4x) Deblur
[…]
Score-SDE 0.33 0.63 94.33 0.58 0.39 53.22 0.63 0.36 66.81
ILVR - - - 0.74 0.28 52.82 - - -
DPS 0.610 0.38 58.89 0.50 0.46 56.08 0.58 0.38 52.64
LGD-MC - 0.16 28.21 - 0.23 34.44 - 0.23 32.57
MPGD 0.75 0.22 11.83 0.79 0.25 60.21 0.77 0.23 51.59
DPS+DSG - 0.12 15.77 - 0.21 30.30 - 0.21 28.22
FMPS-free (Our) 0.92 0.03 8.20 0.77 0.14 24.39 0.76 0.13 20.11
FMPS-gradient (Our) 0.95 0.02 3.29 0.78 0.11 17.83 0.77 0.11 18.49
Table 2: Qualitative evaluation of non-linear inverse problems on CelebA-HQ, where Segmentation is the segmentation maps guided generation, Sketch is the sketch image guided generation, and Face ID is the classifier logit for face image guided generation.
Method Segmentation Sketch Face ID
[…]
FreeDom 1657.0 38.65 34.12 52.18 0.57 0.0452
LGD-MC 2088.5 38.99 49.46 54.47 0.68 0.0445
MPGD 1976.0 39.81 37.23 54.18 0.58 0.0445
FMPS-free (Our) 1427.5 40.36 19.25 60.61 0.46 0.0349
FMPS-gradient (Our) 1213.3 34.10 19.31 52.01 0.39 0.0287
5 Experimental Results
5.1 Implementation Details
Linear inverse problems. Following the previous methods [1], we evaluate FMPS on the CelebA-HQ dataset [26] to test its performance on the linear inverse problems. The linear inverse problems include Inpainting (BOX), Super-resolution (4X), and Gaussian Deblur (kernel size: 61, intensify: 3). The baselines are from the latest methods including Score-SDE [9], ILVR [9], LGD-MC [8], DPS [2], MGPD [3], and DPS+DSG [7]. We use FMPS-free and FMPS-gradient for the linear inverse problem.
Table 3: Ablation study for comparing the FMPS and $\pi$Flow in the linear inverse problem on AFHQ. To make a fair comparison, we choose AFHQ dataset [27] used in $\pi$Flow. We implement $\pi$Flow on the same pre-trained FMs as the FMPS, where the hyperparameters for $\pi$Flow remain in the same setting as reported in its paper. We choose FMPS-gradient since $\pi$Flow relies on the gradient of FMs.
Method SR (4x) Deblur
[…]
$\pi$Flow 0.02 0.64 21.22 0.02 0.63 22.02
FMPS-gradient (Our) 0.69 0.13 7.82 0.72 0.12 7.83
Table 4: Ablation study for comparing the FMPS-gradient and FMPS-free based on the SR (4x) on AFHQ.
[…]
FMPS-gradient 0.74 0.12 7.68 15
FMPS-free 0.69 0.13 7.82 7
Non-linear inverse problem. Following the previous methods, we also evaluate FMPS on the CelebA-HQ dataset to test its performance on the non-linear inverse problems. The non-linear inverse problems include Segmenation maps, Sketch, and Face ID. We use the same pre-trained FMs as the linear inverse problems. The baselines are from the latest methods, including Freedom [1], LGD-MC [8], and MPGD [3]. We use FMPS-free and FMPS-gradient for the non-linear inverse problem.
Text-style Generation. Text-style generation guides the generation using text prompts and the given style image. We first evaluate FMPS on the 1000 text-image pairs. The text prompts are randomly selected from PartiPrompts dataset [1]. Then, the style images are from FreeDom, following the previous works. We use FMPS-free for this task to show the potential of FMPS in large FMs since FMPS-gradient cannot afford the GPU cost for calculating the gradient of Stable Diffusion 3.0.
Metrics. All the distance metrics $D(\ast,\ast)$ we use are the same as previous works, where we have listed more details in the supplementary material. For linear inverse problems, we use structural similarity (SSIM) [28], LPIPS [29], and Fréchet inception distance (FID) [30] as the metric to measure its generation quality. For non-linear inverse problems, we use the norm distance between the condition and generated images, FID, and kernel inception distance (KID) [31] as the metrics. Specifically, the distance norm in Face ID tasks is the classifier logit. For the text-style generation, we use the CLIP score [32] and the Gram matrix [1] to measure the distance between images and text prompts and style images, respectively.
Table 5: Qualitative evaluation of style text-to-image generation, where CLIP [32] is the clip score between the text prompt and the image. Style is the Gram matrix [1] value between generated and style images.
[…]
FreeDom 498.8 30.14
LGD-MC 404.0 21.16
MPGD 441.0 26.61
FMPS-free (Our) 435.2 32.65
Pre-trained Models and settings. To ensure reproduction, we use the pre-trained checkpoint trained in the CelebA-HQ dataset from official Rectified flow [12] for the linear and non-linear inverse problems. For the style task, we use the Stable Diffusion 3 to show the potential of FMPS. The time steps $T$ for all methods in linear and non-linear inverse problems, including FMPS, are set to $T=100$ without specific illustration. All the experiments run on a single RTX 4090 GPU.
Figure 2: The qualitative examples for style tasks using the complex prompts.
5.2 Qualitative Results
Linear inverse problems. To show the ability of the FMPS in the linear inverse problem, we report the comprehensive qualitative evaluation in Table 1. It can be noticed that both FMPS-free and FMPS-gradient generate high-quality images by leveraging the potential of FMs. Concretely, FMPS-free achieves the highest performance on all metrics in Inpainting tasks. Then, in super-resolution tasks, FMPS-free reduces the LPIPS from 0.21 to 0.14 while improving the FID from 30.3 to 24.39. In the end, in the Gaussian Deblur tasks, FMPS-free reduces the LPIPS from 0.21 to 0.13 while improving the FID from 28.22 to 20.11. Compared to the FMPS-free, the FMPS-gradient makes a further improvement since the FMPS-gradient leverages the gradient of the FMs. Concretely, FMPS-gradient improves the 4.91 FID metric in Inpainting tasks. Then, in the super-resolution and Gaussian Deblur tasks, the FMPS-gradient achieves the best performance. These results prove the validity of the FMPS in the linear inverse problems.
Non-linear inverse problems. We further explore the potential of the FMPS in non-linear inverse problems. We report the comprehensive qualitative evaluation in Table 2. It can be noticed that FMPS-free and FMPS-gradient remain in the high-level generation. Concretely, FMPS-gradient achieves state-of-the-art (SOTA) performance in the segmentation task. Meanwhile, in the sketch task, FMPS-gradient improves by an average 43% improvement in the Distance metric while keeping the lowest FID. Meanwhile, FMPS-free greatly improves the distance metric, although the FID has increased. This proves that the gradient information is useful.
Text-style generation. To prove that the FMPS could work on the real generation tasks, we first report the comprehensive qualitative evaluation in Table 5 using 1000 text-image pairs under nine style types. It can be noticed that FMPS-free could better obey the guide for the text prompts since it achieves the highest CLIP score while keeping the style distance at a low level. Then, to show the ability of FMPS-free, we report the qualitative results using the complex text prompts shown in Fig. 2. It can be noticed that FMPS-free achieves a better generation since it maintains a similar style, following the style images while satisfying the text prompts.
5.3 Ablation Study
The ablation study first compares the ability between FMPS and $\pi$Flow. Then, we explore the influence of the $\rho$. In the end, we explore the performance of the FMPS-gradient FMPS-free.
Comparison between FMPS and $\pi$Flow. To show the improvement of FMPS compared to $\pi$Flow, we report the comprehensive results shown in Table 3, where the qualitative examples are shown in Fig. 3. It can be seen that the FMPS-gradient achieved SOTA results. Concretely, FMPS improves the FID from 21.22 to 7.82 and 22.02 to 7.83 in SR (4x) and Deblur tasks, respectively. FMPS also significantly improves SSIM and LPIPS metrics. These results prove the flexibility of the FMPS since it could be implemented in pre-trained FMs and maintain high-quality generation.
Figure 3: The qualitative results for comparing FMPS and $\pi$Flow.
Trade-off between FMPS-gradient and FMPS-free. To illustrate the trade-off between FMPS-gradient and FMPS-free, we report the ablation study in Table 4. The advantage of FMPS-free is the low time cost while it achieves 2x speed-up after dropping the gradient. Meanwhile, its performance has decreased at a low level.
Figure 4: The qualitative results for different $r$. The prompt is “A cat holding a sign that says hello world."
Influence of $r$. To illustrate the influence of $r$, we report the qualitative example based on the text-style generation tasks in Fig. 4. The $r$ is in $[0.3,2]$ from the small to the large. The experimental results show that FMPS works well in different $r$. The difference is the influence of the prompt. In the middle level, i.e., $r\in[0.5,1.5]$, FMPS could ensure that the generated images obey the prompt and change the style correctly.
6 Conclusion
We proposed a novel training-free conditional method for FMs called FMPS, where the key step is explicitly defining the score function in the FMs’ ODEs. Thus, the posterior sampling from DPMs could work on the FMs. We further propose FMPS-gradient and FMPS-free to achieve the posterior sampling. The difference is whether we need to calculate the gradient of the FMs during the conditional generation process, where we offer a trade-off solution to reduce the computation cost. The experimental results show that FMPS could better guide the condition while maintaining high-quality generation.
Limitation. There are some constraints for FMPS, similar to the posterior sampling-based works on DPMs. For example, the seed is the key to the generation since if the seed is bad, the process will fail. Then, the color of the generated images will show a slight deviation compared to the DPM methods, which is an interesting finding since SSIM is not well reported in the linear inverse task. This may be due to the influence between stochastic and arbitrary paths. We will work on these further.